The Gecko Company and the Gordon Company are two firms that have the same business risk but different dividend policies. Gecko pays no dividend, whereas Gordon has an expected dividend yield of 9 percent. Suppose the capital gains tax rate is zero, whereas the income tax rate is 30 percent. Gecko has an expected earnings growth rate of 17 percent annually, and its stock price is expected to grow at this same rate. The aftertax expected returns on the two stocks are equal (because they are in the same risk class).

  

What is the pretax required return on Gordon’s stock?

(Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)
Answer #1

After-tax return = g+D×(1-Tax rate) = 17%

17% = g+9%×(1-30%)

g = 10.7%

Pre-tax return:

= g+D

= 10.7%+9%

= 17.9%

Hence, Pre-tax return is 17.90%

*Please rate Thumbs up